# MACEval: A Multi-Agent Continual Evaluation Network for Large Models

### v1

@@ 3.4 A Framework for MACEval
The proposed MACEval network is configured by three main factors: 1) the role assignment defines the basic functionality of network nodes; 2) the evaluation topology determines the sequential order and interdependencies within the evaluation process; 3) the evaluation network-based metrics and the evaluating continually mechanism provides comprehensive and sustainable capability assessment.
Figure 4: Example of an evaluation network, where colored clusters represent evaluation agents for different capabilities and red arrows denote activated evaluation routes.
@@
Role Assignment. As shown in Fig. 2, the evaluation process for large models is conceptualized as an interview-like procedure, where tasks are posed sequentially or in parallel for probing different capabilities. We mainly involve three types of agents, i.e., the interviewer, the interviewee, and the supervisor. Among them, the interviewee, namely the evaluated model, generates responses according to the given queries. The interviewer performs real-time performance evaluations and determines whether to transform a given question into a new one with dynamically increasing difficulty based on the interviewee’s responses. To ensure the validity of the evaluation process and to avoid scenarios such as repetitive questioning or posing questions that do not meet evaluation requirements, we deploy a third-party model distinct from the evaluation and interviewer models to serve as a supervisory component.
@@
Evaluation Topology. Prior benchmarks run each task independently and average the task metrics on separate subsets of the collected dataset to get an overall metric, neglecting the inherent dependencies between tasks while failing to support progressive and sustainable evaluation (Chang et al., 2024). In MACEval, we introduce four evaluation topologies, including line, star, tree, and hybrid network, to establish the relationships among different evaluation streams. This also enables users to construct customized evaluation networks that better align with the assessment requirements of their target capability. More details are in Appendix C.
@@
${\text{ACC-AUC}}=\int_{a}^{\arg\mathop{\min}\limits_{t}\{t|{\text{ACC}}(t)=0\}}{{\text{ACC}}(t)dt,}\ \ \ {\text{ACC}}(t)=\frac{\sum_{r=1}^{\mathcal{Q}_{total}}\mathbf{1}\left\{\mathcal{M}^{E}\left(q_{r,t}\right)\cong a_{r,t}\right\}}{\mathcal{Q}_{total}},$ (3)
where $a$ is the initial difficulty level, and $\mathbf{1}\left\{\cdot\right\}$ represents the indicator function, which returns 1 if the condition inside the braces holds, and 0 otherwise. $q_{r,t}$} and $a_{r,t}$} are the $r$-th query-answer pair at difficulty level $t$. $Q_{total}$ denotes the total number of query-answer pairs at a certain level. This simple and generally-applicable strategy enables us to evaluate the large models longitudinally and sustainably, as verified in our experimental analysis (Fig. 5).
3.5 Evaluation Routing-based Metric
Building upon the single-dimensional evaluation strategy discussed previously, we introduce a novel evaluation routing-based metric to measure the overall capability of large models. As illustrated in Fig. 4, given an evaluation network with activated evaluation routing $\mathcal{P}$, the comprehensive ability of the interviewee model $\mathcal{M}^{E}$ can be measured as follows:
$\epsilon_{\text{overall}}=\sum_{\left(\mathcal{M}^{E},\mathcal{M}^{I}\right)\in\mathcal{P}}\text{ACC-AUC}_{(\mathcal{M}^{E},\mathcal{M}^{I})}.$ (4)
This metric quantifies the overall model performance as the evaluation network energy derived from edge (evaluation route) weights. It considers the intrinsic dependencies among different capability assessments while enhancing user-oriented customization, thereby enabling a self-organizing and scalable evaluation framework.
Table 2: The performance of different interviewees (LLMs) on 4 text-only capabilities under a 1-hop line evaluation topology configuration. We report ACC-AUC and the difficulty level corresponding to the maximum performance, separated by a slash. The best results are highlighted in bold.
Interviewee (LLM) Interviewer Text Comprehension Math Algorithm Coding
GPT-4o 8.857/10 6.043/7 2.6/6 1.7/2 0.4/1 3.1/4 5.5/10 4.1/6 6.1/11 14.3/16
GPT-4.1 GPT-4o 8.438/9 4.815/6 3.6/10 1.8/2 0.3/1 2.6/4 5.2/10 3.7/6 6.3/12 14.8/16
Gemini 1.5 Pro 8.766/10 5.289/6 1.4/3 1.3/2 0.5/1 2.8/3 5.4/10 4.2/6 6.1/12 14.6/18
Gemini 2.0 Flash 8.898/10 5.049/6 2.8/8 1.6/2 0.6/1 3.1/4 6.1/10 4.5/7 7.3/13 16.6/22
Gemini 2.5 Pro GPT-4o 8.978/10 6.062/7 3.8/9 3.3/5 0.9/3 3.2/4 6.2/10 4.9/8 7.8/15 17.8/22
DeepSeek-V3 5.293/6 5.099/6 2.0/5 2.6/4 0.7/1 3.0/3 4.4/10 4.6/6 6.4/13 15.3/17
DeepSeek-R1 GPT-4o 9.871/10 5.936/7 8.3/10 7.2/10 0.9/1 3.1/5 9.3/13 7.2/9 7.5/13 17.9/20
Qwen3-8B 7.626/10 2.846/4 1.4/4 1.0/1 0.9/2 1.1/2 4.2/8 2.8/4 5.8/8 11.5/14
Qwen2.5-7B 6.797/8 3.003/4 0.9/2 0.9/1 0.5/1 1.6/2 3.4/8 2.6/4 5.2/8 9.8/13
Qwen2.5-14B 7.903/10 3.327/4 2.0/5 1.0/1 0.5/1 2.2/3 4.4/8 3.0/4 5.5/8 11.2/13
Qwen2.5-72B 8.543/10 4.274/5 2.2/5 1.0/1 0.5/1 2.4/3 2.8/4 3.7/5 5.8/9 13.9/15
Qwen2-7B GPT-4o 7.339/9 3.053/4 0.5/1 0.7/1 0.4/1 1.2/2 2.8/7 2.4/4 4.7/7 8.5/13
Llama3.3-70B 7.619/10 4.564/6 2.8/8 1.0/1 0.5/1 2.6/3 4.1/8 3.3/4 5.7/9 13.6/14
Llama3.2-3B 5.641/8 2.369/3 0.1/1 0.3/1 0.3/1 0.6/1 1.4/4 1.7/3 2.2/5 7.3/8
Llama3.1-8B GPT-4o 6.740/9 3.006/4 0.7/4 0.2/1 0.4/1 0/0 2.9/8 2.1/4 4.6/7 9.3/12
@@ 4.1 Experimental Setup
Benchmark Candidates. Our experiments include 23 large models total, with a mix of cutting-edge proprietary models, open-source LLMs, and MLLMs. Specifically, for proprietary models, we include OpenAI models such as GPT-4o (OpenAI, 2024), and GPT-4.1 (2025-04-14) (Achiam et al., 2023), Google models such as Gemini-1.5-Pro, Gemini-2.0-Flash (Reid et al., 2024), and Gemini-2.5-Pro (Kavukcuoglu, 2025). For open-source LLMs, we include three mainstream language backbones widely used in many large models, i.e., the DeepSeeks (DeepSeek-V3 (Liu et al., 2024a) and DeepSeek-R1 (Guo et al., 2025)), the Qwens (Qwen3-8B (Team, 2025), Qwen2.5-{7, 14, 72}B (Team, 2024), Qwen2-7B (Yang et al., 2024a)), and the Llamas (Llama3.3-70B, Llama3.2-3B, Llama3.1-8B) (Grattafiori et al., 2024). For open-source MLLMs, we include models such as Qwen2.5-vl-{3, 7, 72}B (Bai et al., 2025), Qwen2-vl-7B (Wang et al., 2024), InternVL3-8B (Zhu et al., 2025), InternVL2.5-{8, 38}B (Chen et al., 2024d), InternVL2-8B (Chen et al., 2024e).
Figure 5: Performance curves and ACC-AUC values of different series of MLLMs. The positions of the red circles indicate the upper bound of the model’s capabilities. The upper and lower parts of the subfigure depict the IQP and CU tasks, respectively.
Implementation Details. For all models, we perform evaluation using their respective templates under zero shot settings. The initial difficulty level $a$ and the total number of queries in each level $\mathcal{Q}_{total}$ are set to 1 and 10, respectively. The max_new_tokens is set to the maximum value supported by each model to ensure the completeness of the response. We employ SSH connections to establish communication between the local interviewer model and the remote interviewee model running on a server. All experiments are conducted using a maximum of 8 RTX4090 24GB GPUs.
@@ 4.2 Main Results
In Tab. 2 and Fig. 5, we analyze the performance of 15 LLMs and 8 MLLMs under different settings on 9 open-ended tasks. Our evaluation brings several important findings, as follows:
@@
4) Steep decline in low-level visual perception. In Fig. 5, we observe a sharp performance cliff in the low-level IQP tasks, which suggests potential Just Noticeable Difference (JND) points in the vision backbones of MLLMs. Besides, the InternVL series exhibits more unstable perceptual responses compared to the Qwens and other proprietary models.
Figure 7: Maximum performance level vs. ACC-AUC. We plot the performance of 15 LLMs compared to the corresponding maximum difficulty levels with each a fit line and correlation coefficient. The red dotted lines denote the performance saturation line.
4.3 Correlation Analyses
In Fig. 6, we compute the Pearson correlation coefficient among all pairs of tasks. Results show that, unsurprisingly, text comprehension, algorithm reasoning, and coding all correlate with one another (avg. $r>0.81$). Moreover, math tasks involving intensive numerical computation correlate relatively weakly with all other categories, which, together with the extremely low maximum capability level reported in Tab. 2, indicates the uniqueness of this capability. We further investigate the relationship between maximum capability level and long-term ACC-AUC in Fig. 7. It can be observed that these two dimensions are highly correlated (avg. $r\approx 0.888$) in all tasks, and few models reach saturated performance within a limited capability level.
1-hop line (w/o) 36% 42% 18% 2% 50% 52% 26% 36% 30% 18%
Node sequence STUdisrupt.->STUmask. Arith.scale->Arith.oper. SPSnode->SPSedge Generation->Debugging
2-hop line (w/) 18% (-24%) 4% (-46%) 22% (-14%) 14% (-4%)
@@ 4.4 Effect of Different Evaluation Topologies
1) Message passing improves the quality of question generation. We focus on two types of question generation errors caused by hallucinations from the interviewer model during the evaluation process, i.e., format violations and redundant questions. Tab. 3 demonstrates the effectiveness of the message passing mechanism in improving the question quality between two related tasks with an average error rate reduction of 22%, which serves as a special prompt engineering. Since the subject of the visual perception tasks is the image itself, no obvious mistakes in question generation were detected.
Table 4: Results of different interviewers.
@@
2) Overall performance comparison based on the evaluation network energy. As shown in Fig. 10, we compute the overall performance for 15 LLMs based on the evaluation routes in Fig. 4. To ensure fairness, we remove two 1-hop visual perception-related evaluation nodes, which are not affected by the message-passing mechanism. We notice that proprietary models still maintain a dominant position. The characteristic of $\epsilon_{\text{overall}}$ lies in considering underlying inter-task dependencies, incorporating information-aware evaluation routing instead of independently assessing and aggregating task scores.
4.5 Analysis of Interviewer Diversity
Next, we discuss the performance differences under different interviewer settings. As listed in Tab. 4, the evaluation results remain consistent across different interviewers in most scenarios. However, when evaluating tasks that involve more complex text generation, using a weaker model as the interviewer may limit the exploration of the interviewee model’s upper-bound capabilities. See more experimental results in the Appendix. D.

### v2

@@ 3.3 MACEval Framework
As shown in Fig. 2, the proposed MACEval network is configured by three main factors: 1) the role assignment defines the basic functionality of network nodes; 2) the evaluation topology determines the sequential order and interdependencies within the evaluation process; 3) the evaluation network-based metrics and the evaluating continually mechanism provides comprehensive and sustainable capability assessment.
Table 1: Performance comparison on four text-only capabilities under a 1-hop line evaluation topology. We report the ACC-AUC/Maximum performance, as well as the utility-capability ratio. The best results are highlighted in bold.
Interviewee (LLM) Interviewer Text Comprehension Math Algorithm Coding $\epsilon_{\text{overall}}$
GPT-4o GPT-4o 8.857/10 6.043/7 2.6/6 1.7/2 0.4/1 3.1/4 5.5/10 4.1/6 6.1/11 14.3/16 6.889
GPT-4.1 8.438/9 4.815/6 3.6/10 1.8/2 0.3/1 2.6/4 5.2/10 3.7/6 6.3/12 14.8/16 6.537
Gemini 1.5 Pro 8.766/10 5.289/6 1.4/3 1.3/2 0.5/1 2.8/3 5.4/10 4.2/6 6.1/12 14.6/18 6.868
Gemini 2.0 Flash 8.898/10 5.049/6 2.8/8 1.6/2 0.6/1 3.1/4 6.1/10 4.5/7 7.3/13 16.6/22 6.825
Gemini 2.5 Pro 8.978/10 6.062/7 3.8/9 3.3/5 0.9/3 3.2/4 6.2/10 4.9/8 7.8/15 17.8/22 6.508
DeepSeek-V3 5.293/6 5.099/6 2.0/5 2.6/4 0.7/1 3.0/3 4.4/10 4.6/6 6.4/13 15.3/17 7.081
DeepSeek-R1 9.871/10 5.936/7 8.3/10 7.2/10 0.9/1 3.1/5 9.3/13 7.2/9 7.5/13 17.9/20 7.892
Qwen3-8B GPT-4o 7.626/10 2.846/4 1.4/4 1.0/1 0.9/2 1.1/2 4.2/8 2.8/4 5.8/8 11.5/14 6.596
Qwen2.5-7B 6.797/8 3.003/4 0.9/2 0.9/1 0.5/1 1.6/2 3.4/8 2.6/4 5.2/8 9.8/13 6.729
Qwen2.5-14B 7.903/10 3.327/4 2.0/5 1.0/1 0.5/1 2.2/3 4.4/8 3.0/4 5.5/8 11.2/13 7.104
Qwen2.5-72B 8.543/10 4.274/5 2.2/5 1.0/1 0.5/1 2.4/3 2.8/4 3.7/5 5.8/9 13.9/15 7.460
Qwen2-7B 7.339/9 3.053/4 0.5/1 0.7/1 0.4/1 1.2/2 2.8/7 2.4/4 4.7/7 8.5/13 6.104
Llama3.3-70B 7.619/10 4.564/6 2.8/8 1.0/1 0.5/1 2.6/3 4.1/8 3.3/4 5.7/9 13.6/14 7.181
Llama3.2-3B 5.641/8 2.369/3 0.1/1 0.3/1 0.3/1 0.6/1 1.4/4 1.7/3 2.2/5 7.3/8 5.064
Llama3.1-8B 6.740/9 3.006/4 0.7/4 0.2/1 0.4/1 0/0 2.9/8 2.1/4 4.6/7 9.3/12 4.595
@@
Role Assignment. We mainly involve three types of agents, i.e., the interviewer, the interviewee, and the supervisor. Among them, the interviewee, namely the evaluated model, generates responses according to the given queries. The interviewer performs real-time performance evaluations and determines whether to transform a given question into a new one with dynamically increasing difficulty based on the interviewee’s responses. To ensure the validity of the evaluation process and to avoid scenarios such as repetitive questioning or posing questions that do not meet evaluation requirements, we deploy a third-party model distinct from the evaluation and interviewer models to serve as a supervisory component.
@@
Evaluation Topology. Prior benchmarks run each task independently and average the task metrics on separate subsets of the collected dataset to get an overall metric, neglecting the inherent dependencies between tasks while failing to support progressive and sustainable evaluation (Chang et al., 2024). In MACEval, we introduce four evaluation topologies, including line, star, tree, and hybrid networks, to establish the relationships among different evaluation streams. This also enables users to construct customized evaluation networks that better align with the assessment requirements of their target capability.
Figure 4: Performance curves and ACC-AUC values of different MLLM series. The red circles indicate the upper bound of the model’s capabilities. The left and right parts in subfigures depict the IQP and CU tasks, respectively.
@@
ACC-AUC $=\int_{a}^{\arg\mathop{\min}\limits_{t}\{t|{\text{ACC}}(t)=0\}}{{\text{ACC}}(t)dt,}$ (3)
${\text{ACC}}(t)$ $=\frac{\sum_{r=1}^{\mathcal{Q}_{total}}\mathbf{1}\left\{\mathcal{M}^{E}\left(q_{r,t}\right)\cong a_{r,t}\right\}}{\mathcal{Q}_{total}},$
where $a$ is the initial difficulty level, and $\mathbf{1}\left\{\cdot\right\}$ represents the indicator function, which returns 1 if the condition inside the braces holds, and 0 otherwise. $q_{r,t}$} and $a_{r,t}$} are the $r$-th query-answer pair at difficulty level $t$. $Q_{total}$ denotes the total number of query-answer pairs at a certain level. This simple and generally-applicable strategy enables us to evaluate the large models longitudinally and sustainably, as verified in our experimental analysis.
3.4 Utility-Capability Ratio
Building upon the single-dimensional evaluation strategy discussed previously, we introduce an evaluation routing-based metric to measure the overall utility-capability ratio of large models. Given an evaluation network with activated evaluation routing $\mathcal{P}$, the overall utility-capability ratio of the interviewee model $\mathcal{M}^{E}$ can be computed as follows:
$\epsilon_{\text{overall}}=\sum_{\left(\mathcal{M}^{E},\mathcal{M}^{I}\right)\in\mathcal{P}}\frac{\text{ACC-AUC}_{(\mathcal{M}^{E},\mathcal{M}^{I})}}{\arg\min_{t}\left\{t|\text{ACC}\left(t\right)=0\right\}}$ (4)
which quantifies the overall performance utilization under the limited capacity.
@@ 4.1 Experimental Setup
Benchmark Candidates. Our experiments include 23 large models total, with a mix of cutting-edge proprietary models, open-source LLMs, and MLLMs. Specifically, for proprietary models, we include OpenAI models such as GPT-4o (OpenAI, 2024), and GPT-4.1 (2025-04-14) (Achiam et al., 2023), Google models such as Gemini-1.5-Pro, Gemini-2.0-Flash (Reid et al., 2024), and Gemini-2.5-Pro (Kavukcuoglu, 2025). For open-source LLMs, we include three mainstream language backbones widely used in many large models, i.e., the DeepSeeks (DeepSeek-V3 (Liu et al., 2024a) and DeepSeek-R1 (Guo et al., 2025)), the Qwens (Qwen3-8B (Team, 2025), Qwen2.5-{7, 14, 72}B (Team, 2024), Qwen2-7B (Yang et al., 2024a)), and the Llamas (Llama3.3-70B, Llama3.2-3B, Llama3.1-8B) (Grattafiori et al., 2024). For open-source MLLMs, we include models such as Qwen2.5-vl-{3, 7, 72}B (Bai et al., 2025), Qwen2-vl-7B (Wang et al., 2024), InternVL3-8B (Zhu et al., 2025), InternVL2.5-{8, 38}B (Chen et al., 2024b), InternVL2-8B (Chen et al., 2024c).
Figure 5: Utility-capability ratio comparison of five representative models across different dimensions.
Implementation Details. For all models, we perform evaluation using their respective templates under zero-shot settings. The initial difficulty level $a$ and the total number of queries in each level $\mathcal{Q}_{total}$ are set to 1 and 10, respectively. The max_new_tokens is set to the maximum value supported by each model to ensure the completeness of the response. We employ SSH connections to establish communication between the local interviewer model and the remote interviewee model running on a server. All experiments are conducted using a maximum of 8 RTX4090 24GB GPUs.
@@ 4.2 Main Results
In Tab. 1, Fig. 4, and Fig. 5, we analyze the performance of 15 LLMs and 8 MLLMs under different settings on 10 open-ended tasks. Our evaluation brings several important findings, as follows:
@@
Figure 6: Pearson correlation between different tasks.
Figure 7: Maximum performance level vs. ACC-AUC. We plot the performance of 15 LLMs compared to the corresponding maximum difficulty levels, with each a fit line and correlation coefficient. The red dotted lines denote the performance saturation line.
@@
4) Performance utilization varies across models and tasks. As shown in Tab. 1, DeepSeek-R1 has the highest overall utility-capability ratio. While the Qwen2.5 series surpasses its next generation Qwen3-8B, indicating the possible instability resulting from the combination of the reasoning mode. Fig. 5 exhibits the discrepancies in performance utilization among tasks, where most models show low correlation between their performance utilization rate and absolute performance ranks, suggesting another optimization direction for large models.
5) Steep decline in low-level visual perception. In Fig. 4, we observe a sharp performance cliff in the low-level IQP tasks, which suggests potential Just Noticeable Difference (JND) points in the vision backbones of MLLMs. Besides, the InternVL series exhibits more sensitive perceptual responses compared to the Qwens and other models.
4.3 Correlation Analysis
In Fig. 6, we compute the Pearson correlation coefficient among all pairs of tasks. Results show that, unsurprisingly, text comprehension, algorithm reasoning, and coding all correlate with one another (avg. $r>0.81$). Moreover, math tasks involving intensive numerical computation correlate relatively weakly with all other categories, which, together with the extremely low maximum capability level reported in Tab. 1, indicates the uniqueness of this capability. We further investigate the relationship between maximum capability level and long-term ACC-AUC in Fig. 7, where two dimensions are highly correlated (avg. $r\approx 0.888$) in all tasks, and few models reach saturated performance within a limited capability level.
@@ 4.4 Effect of Different Evaluation Topologies
1) Ablation Study on Interviewer Diversity. Here, we discuss the performance differences under different interviewer settings. As listed in Tab. 2, the evaluation results remain consistent across different interviewers in most scenarios. However, when evaluating tasks that involve more complex text generation, using a weaker model as the interviewer may limit the exploration of the interviewee model’s upper-bound capabilities.
Table 2: Ablation study on different interviewers.
@@
Table 3: Results on different role settings. Triple means Qwen2.5-7B, GPT-4o, and Claude 3.7 Sonnet are employed as interviewers.
Interviewee Interviewer Supervisor STUdisrupt. Arith.scale BTT
Qwen2.5-7B GPT-4o Claude 3.7 Sonnet 6.797/8 0.9/1 1.6/2
Qwen2.5-7B Qwen2.5-7B 6.833/8 1.0/2 1.6/2
Triple Qwen2.5-7B 6.802/8 0.9/1 1.6/2
GPT-4o GPT-4o Claude 3.7 Sonnet 8.855/10 1.7/2 3.1/4
GPT-4o GPT-4o 8.855/10 1.7/2 3.1/4
1-hop line (w/o) 36% 42% 18% 2% 50% 52% 26% 36% 30% 18%
[…]
2-hop line (w/) 18% (-24%) 4% (-46%) 22% (-14%) 14% (-4%)
2) Multiple Roles Exploration. We further evaluate the effect of different role settings, where the same model is used simultaneously as the interviewee, interviewer, and supervisor. Concretely, we test Qwen2.5-7B and GPT-4o in three tasks, i.e., STUdisrupt., Arith.scale, and BTT. In Tab. 3, we observe that changing the supervisor does not explicitly affect the resulting performance. Meanwhile, after changing the interviewer from GPT-4o to Qwen2.5-7B, the evaluated performance of Qwen2.5-7B on STUdisrupt. and Arith.scale tasks grows slightly. We suspect that this is due to the limited diversity of the generated queries stemming from the relatively weak capability of the interviewer model (similar observation as Tab. 2). Therefore, we employ a 1v3 star evaluation topology to solve the mild variance for the same interviewee model across interviewers, which indeed improves the reliability of the evaluation process. Moreover, we find a high error rate (56%) in the BTT task, calculated by the misalignment between the external function and self-generated ground-truth, when Qwen2.5-7B serves as the interviewer. This result is much lower in GPT-4o’s setting, showing the necessity of using relatively powerful models and incorporating external functions.
@@
3) Overall performance comparison based on the evaluation routing. In Fig. 13, we show the sum of normalized performance for 15 LLMs based on the given evaluation routes. We notice that proprietary models still maintain a dominant position. The characteristic of $\epsilon_{\text{overall}}$ lies in considering underlying inter-task dependencies, incorporating information-aware evaluation routing instead of independently assessing and aggregating task scores. We then compare these results to other mainstream benchmarks (LiveBench (White et al., 2025) and ChatBot Arena (Chiang et al., 2024)), shown in Fig. 9. It can be observed that the results of MACEval highly correlate with mainstream benchmarks (avg. $r=0.9509$), demonstrating the effectiveness of MACEval and the scalable design concept of MACEval, where models can conduct fair performance tests within a fixed evaluation network.
Figure 8: Error rate (%) versus propagation round.
4.5 Potential Error Propagating
1) Message passing improves the quality of question generation. We focus on two types of question generation errors caused by hallucinations from the interviewer model during the evaluation process, i.e., format violations and redundant questions. Tab. 4 demonstrates the effectiveness of the message passing mechanism in improving the question quality between two related tasks with an average error rate reduction of 22%. For MLLM evaluations, since the subject of the visual perception tasks is the image itself, no obvious mistakes in question generation were detected.
2) Impact of Proactive Error Injection. We conduct extra experiments to explore the robustness of MACEval framework. Specifically, we add interference characters (unordered and meaningless strings) into the message, which are then taken to the next node. We evaluate the data generation process of the binary tree traversal (BTT) task, which has the highest initial error (52%, Tab. 4). The results in Fig. 8 show that injecting irrelevant textual contents exhibits little influence on the inherent error rate that depends on the difficulty of the task itself. Conversely, the progressive accumulation of erroneous content renders the information more readily detectable by a third-party supervisor, facilitating its timely suppression and the issuance of a regeneration directive to the interviewer model.
Task\#Query 10 20 30 50
STUmask. 4.247/5 4.256/5 4.253/5 4.244/5
BTT 2.40/3 2.50/3 2.43/3 2.46/3
Table 6: Results of the statistical test.
#Query@N times STUmask. BTT
10@10 [4.2444, 4.2558] [2.4122, 2.4618]
20@10 [4.2457, 4.2533] [2.4265, 2.4570]
pp-value 0.8486 0.7227
4.6 Stability Analysis
To validate the stability of evaluation data generation, we compare the performance differences in terms of the number of generated queries at the same difficulty level on STUmask. and BTT tasks, as shown in Tab. 5. The interviewee and interviewer are Qwen2.5-72B and GPT-4o, respectively. We also conduct two-sided t-test and reported 95% PI and $p$-value for 300 queries in Tab. 6. We can observe that there is no significant difference between the results of 10@10 and 20@10, as the $p$-value>>0.05. Their 95% PIs have a considerable overlap, indicating the robustness of the evaluation process. Furthermore, the 95% PI of 20@10 is more concentrated than that of 10@10, showing the precision of increasing the number of rounds under the given difficulty level. See more experimental results in the Appendix. C.
4.7 Comparison to Other Large Model Benchmarks
We compare the normalized evaluation results obtained by MACEval to popular benchmarks (LiveBench (White et al., 2025) and ChatBot Arena (Chiang et al., 2024)), shown in Fig. 9. It can be observed that the results of MACEval highly correlate with mainstream benchmarks (avg. $r=0.9509$), demonstrating the effectiveness of MACEval and proving the potential of in-process problem generation.